OpenStudy (ahrimain):

Kayla needs $560.00 to buy a new laptop she already has $200.00 in the bank. Each month, she adds $45.00 to her bank account. In how many more months will she have enough money to buy the laptop A: 5 B: 6 C: 7 D: 8

OpenStudy (anonymous):

D x=months 200+45x=560 -200 -200 45x=360 divide both sides by 45 so 360 divided by 45 is 8 so x=8 then plug in and check 200+45(8)=560